# Moving sum procedure for multiple change point detection in large factor models

### v1

@@ Abstract
The paper proposes a moving sum methodology for detecting multiple change points in high-dimensional time series under a factor model, where changes are attributed to those in loadings as well as emergence or disappearance of factors. We establish the asymptotic null distribution of the proposed test for family-wise error control, and show the consistency of the procedure for multiple change point estimation. Simulation studies and an application to a large dataset of volatilities demonstrate the competitive performance of the proposed method.
11footnotetext: Department of Economics, Università di Bologna. Email: matteo.barigozzi@unibo.it.22footnotetext: School of Mathematics, University of Bristol. Email: haeran.cho@bristol.ac.uk.33footnotetext: Department of Economics and Management, Universita’ di Pavia and School of Business, University of Leicester. Email: lt285@leicester.ac.uk.
@@ 1 Introduction
Factor models are, arguably, one of the most frequently employed tools to model and carry out inference when large-dimensional, vector-valued time series are available. Whilst a comprehensive review of the literature goes well above and beyond the scope of this paper, it is worth noting that factor models have a long history, dating back at least to the seminal paper by Spearman, (1904). Since being popularised by the contribution by Chamberlain and Rothschild, (1983), and since the development of the asymptotic theory to analyse large dimensional factor models (Bai,, 2003) their usage has become de rigueur in social sciences and economics where they have been applied to diverse fields such as business cycle analysis, asset pricing and economic monitoring and forecasting (see, inter alia, the review by Stock and Watson, (2011) for a comprehensive list of references).
@@
When factor models are concerned, on the one hand, the analysis in Stock and Watson, (2009) and Bates et al., (2013) shows that, when such changes are “sufficiently small” and do not involve a change in the number of common factors, the presence of change points is inconsequential for inference on the factor spaces. On the other hand, in several applications, there is no guarantee that the factor structure only undergoes small changes and that the number of common factors is time-invariant. Hence, it is not surprising that the literature has developed several methods to test, retrospectively, whether there is a change point or not. Breitung and Eickmeier, (2011), Corradi and Swanson, (2014), Han and Inoue, (2015), Yamamoto and Tanaka, (2015), Baltagi et al., (2017), Cheng et al., (2016), Duan et al., (2023), and, recently, Bai et al., (2024) propose several tests to check if there is a break, based on the idea that a change point in the loadings is equivalent to a change in the covariance matrix of the common factors. Thus assuming homoscedastic factors, a change in the covariance matrix of the estimated common factors can only be due to a break in the loadings. See also Massacci, (2023) and, in an online context, Barigozzi and Trapani, (2020) and He et al., (2024) for alternative formulation of change point tests.
Whilst, on account of the references above, the literature on testing for change points seems well-established, there is relatively little work when the estimation of a change point is concerned. In particular, the literature largely focuses on the At-Most-One-Change alternative (Duan et al.,, 2023), and few addresses the detection and estimation of (possibly) multiple change points in the factor structure, with some exceptions (Barigozzi et al.,, 2018; Li et al.,, 2023).
@@
In this paper, we make an advance on the change point literature by proposing a method to estimate the number and locations of (possibly) multiple change points under a (static) factor model. Specifically, we propose a procedure based on the so-called MOving SUMs (MOSUM) process which, since the seminal contribution by Hušková and Slabỳ, (2001), has been shown to have desirable properties in multiple change point estimation under weak assumptions on the distribution of the data. Versatility of the MOSUM procedure has been demonstrated in the context of detecting changes in the mean (Hušková and Slabỳ,, 2001; Eichinger and Kirch,, 2018), trend (Kim et al.,, 2024) and the drift of stochastic processes (Kirch and Klein,, 2024); see also Kirch and Reckruehm, (2024) for a general framework for multivariate time series segmentation based on an estimating function. There also exist recent contributions which extend the use of this methodology to high-dimensional time series (Cho and Owens,, 2024; Cho et al.,, 2024), demonstrating its scalability and suggesting that it is worth exploring its performance in the context of large factor models.
@@
We build on that, as discussed above, a change in the loadings is observationally equivalent to a change in the covariance matrix of the common factors, and construct a MOSUM statistic based on the (moving) partial sums of the outer products of the estimated factors. In this respect, our statistic is akin to the one based on the (maximally selected) sequence of likelihood ratio tests proposed e.g. in Duan et al., (2023), save for the fact that it is based on using moving, as opposed to cumulative, partial sums. Upon testing for changes in the factor model using the MOSUM process, if the null hypothesis of no changes is rejected, we perform multiple change point detection by estimating both the total number and the locations of the breaks. We establish the asymptotic null distribution of the MOSUM process as well as consistency in multiple change point detection, with the accompanying rate of estimation for individual breaks. In doing so, we derive a Strong Invariance Principle for the MOSUM process based on the outer products of the estimated factors, which may be of independent interest.
@@
For a random variable $X$, we have $|X|_{\nu}=(\mathsf{E}(|X|^{\nu}))^{1/\nu}$. We define: weak convergence on the space $\mathcal{D}[0,1]$ with $\overset{\mathcal{D}}{\rightarrow}$; =$\overset{\mathcal{D}}{=}$ is equality in distribution; $\rightarrow$ is the ordinary limit; $\Gamma(\cdot)$ is the Gamma function; $\log(x)$ is the natural logarithm of $x>0$; $\mathbb{I}_{\mathcal{A}}$ is an indicator function satisfying $\mathbb{I}_{\mathcal{A}}=1$ if the event $\mathcal{A}$ is true and $\mathbb{I}_{\mathcal{A}}=0$ otherwise. By $\mathbf{I}$, $\mathbf{O}$ and $\mathbf{0}$, we denote an identity matrix, a matrix of zeros and a vector of zeros whose dimensions depend on the context. For a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, we denote by $\mathbf{A}^{\top}$ its transpose and $\|\mathbf{A}\|$ its Euclidean norm, with $\Lambda_{\max}(\mathbf{A})$ and $\Lambda_{\min}(\mathbf{A})$ denoting its largest and the smallest eigenvalues in modulus. When $m=n$, we denote by $\mathsf{Vech}(\mathbf{A})$ the vector of length $m(m+1)/2$ that stacks the elements on and below the main diagonal of $\mathbf{A}$, and $\mathbf{L}_{m}$ the $m(m+1)/2\times m^{2}$-matrix satisfying $\mathsf{Vech}(\mathbf{A})=\mathbf{L}_{m}\mathsf{Vec}(\mathbf{A})$, where $\mathsf{Vec}(\cdot)$ is an operator that stacks the columns of the matrix into a vector. Conversely, we define $\mathbf{K}_{m}\in\mathbb{R}^{m^{2}\times m(m+1)/2}$ satisfying $\mathsf{Vec}(\mathbf{A})=\mathbf{K}_{m}\mathsf{Vech}(\mathbf{A})$, see Appendix A.12 of Lütkepohl, (2005). Finally, given two sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}=O(b_{n})$ if, for some finite positive constant $C$ there exists $N\in\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$ such that $|a_{n}||b_{n}|^{-1}\leq C$ for all $n\geq N$.
@@ 2.1 Model
Here, the matrix of loadings $\bm{\Lambda}_{j}=[\bm{\lambda}_{j,1},\ldots,\bm{\lambda}_{j,N}]^{\top}\in\mathbb{R}^{N\times r}$ has column rank $r_{j}\leq r$ fixed for all $N$, and “loads” the vector of random factors $\mathbf{f}_{t}=(f_{1,t},\ldots,f_{r,t})^{\top}$ onto the cross-sections of $\mathbf{X}_{t}$, and $\mathbf{e}_{t}=(e_{1,t},\ldots,e_{N,t})^{\top}$ denotes the idiosyncratic component. The model in (1) allows for changes due to emergence or disappearance of factor(s) by permitting the ranks of $\bm{\Lambda}_{j}$ to vary, as well as rotational changes in the loading matrices. At the same time, the model is not identifiable in that the changes in the loadings are (observationally) equivalent to changes in the second-order properties of the common factor $\mathbf{f}_{t}$ – a fact, as mentioned in Introduction, frequently explored in the relevant literature for developing change point tests in the At-Most-One-Change scenario; see e.g. Bai et al., (2024). We further illustrate this point in the following example.
@@
Consider the factor model in (1) with $R=1$ and the single change point at $k_{1}=k^{\ast}$. For any $\bm{\Lambda}_{j}\in\mathbb{R}^{N\times r_{j}}$, $j=0,1$, we can find $\bm{\Lambda}\in\mathbb{R}^{N\times r}$ of full column rank with $r\geq\max(r_{0},r_{1})$ such that $\bm{\Lambda}_{j}=\bm{\Lambda}\mathbf{A}_{j}$ for $\mathbf{A}_{j}\in\mathbb{R}^{r\times r}$ of rank $r_{j}$. Then, we can re-write (1) as
@@
$\mathbf{X}_{t}=\bm{\Lambda}\sum_{j=0}^{R}\mathbf{A}_{j}\mathbf{f}_{t}\cdot\mathbb{I}_{\{k_{j}<t\leq k_{j+1}\}}+\mathbf{e}_{t}=\bm{\Lambda}\mathbf{g}_{t}+\mathbf{e}_{t}$ (2)
@@
for $1\leq j\leq R$. Then, the problem of detecting the multiple change points in the loadings under (2), becomes that of detecting change points in the covariance of the pseudo factors $\mathbf{g}_{t}$.
@@ 2.2 Methodology
Let us suppose that the number of pseudo factors $r$ is known. Then, we can estimate the pseudo factors $\mathbf{g}_{t}$ as, up to a rotation, $\sqrt{T}$ times the $r$ leading eigenvectors of the $T\times T$ matrix $(NT)^{-1}\mathbf{X}\mathbf{X}^{\top}$ with $\mathbf{X}=\mathbf{X}_{0:T}$. Denoting such estimator by $\widehat{\mathbf{g}}_{t}$, we propose to test for a change under the model in (1) and, if any, estimate the multiple change points, by scanning the data using the following MOSUM statistic:
@@
For the implementation of the MOSUM procedure, we require an estimator of the long-run covariance matrix $\mathbf{V}_{k}$. While we allow it to be location-dependent to account for the heteroscedasticity in the presence of change points, the difficulty in estimating a long-run covariance matrix of multivariate time series is well-known. In the current setting, this is further augmented by that the computation of $\mathcal{T}_{N,T,\gamma}(k)$ calls for the inverse of $\mathbf{V}_{k}$, which may bring further numerical instabilities. Therefore, we opt to use the following HAC-type estimator in place of $\mathbf{V}_{k}$,
@@
with some bandwidth $m\geq 1$. We later show that the matrix $\widehat{\mathbf{V}}$ provides a consistent estimator of the long-run covariance matrix of $\mathbf{M}_{N,T,\gamma}(k)$ under $\mathcal{H}_{0}$ (see Proposition 2). We note that the use of the Bartlett kernel in (5) is only one of the many possible choices, and we refer to Bai et al., (2024) for a comprehensive analysis of various kernel-based estimators of $\mathbf{V}$. For consistency of multiple change point detection, we only require that a positive definite matrix is used in place of $\mathbf{V}_{k}$ (Theorem 3). By default, we propose to adopt $\widehat{\mathbf{V}}$ (or its diagonal entries) which is shown to work well; see Section 4 for further discussions.
@@ 3.1 Assumptions
Since the seminal works by Wu, (2005) and Berkes et al., (2011) (see also Hörmann,, 2009), decomposable Bernoulli shifts have proven a very convenient way to model and study dependent time series, mainly due to their generality and since it is much easier to verify whether a sequence forms a decomposable Bernoulli shift than, e.g. verifying mixing conditions. Virtually all the most commonly employed models in econometrics and statistics can be shown to generate decomposable Bernoulli shifts, such as ARMA and (G)ARCH processes, non-linear time series models (e.g. random coefficient autoregressive models and threshold models), Volterra series and data generated by dynamical systems; see Berkes et al., (2011), Aue et al., (2009) and Liu and Lin, (2009).
@@
    we have $\mathbf{D}$ invertible.
@@ 3.2 Asymptotic null distribution
The limiting law in Theorem 1 is analogous to those derived in Hušková and Slabỳ, (2001) and Eichinger and Kirch, (2018), modulo the fact that here, we deal with $d$-variate vectors and therefore the function $b_{d}(\cdot)$ depends on $d$. In contrast with a “standard” multivariate time series application, however, in our result, the cross-sectional dimension $N$ also plays a role through the definition of $\zeta$ in (7). As a by-product, we establish the SIP of the process $\mathsf{Vech}(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top})$ after an appropriate centering (see Lemma A.11 in Appendix A.1), which proves crucial in deriving the asymptotic null distribution in (10).
@@
The condition in (8) places both upper and lower bounds on the bandwidth $\gamma$. Specifically, $\gamma$ is required to grow sufficiently faster than $T^{2\zeta}$ while satisfying $T^{-1}\gamma\rightarrow 0$, and the former restriction calls for larger $\gamma$ when $\mathbf{f}_{t}$ has fewer moments or when $N$ is small. We note that $\beta$ is known and does not need to be estimated. Conversely, $\nu$ is in general not known. We may select $\gamma$ satisfying (8) by plugging in an estimate of $\nu$, say $\widehat{\nu}$; alternatively, one can decide a value of $\nu$, say $\nu^{\ast}$, and test whether $\left|\mathbf{g}_{t}\right|_{\nu^{\ast}}<\infty$. In both cases, one difficulty is that $\mathbf{g}_{t}$ is not observable; however, deriving $\nu$ from the data $\mathbf{X}_{t}$ yields a lower bound.

### v2

@@ Abstract
This paper proposes a moving sum methodology for detecting multiple change points in high-dimensional time series under a factor model, where changes are attributed to those in loadings as well as emergence or disappearance of factors. We establish the asymptotic null distribution of the proposed test for family-wise error control, and show the consistency of the procedure for multiple change point estimation. Simulation studies and an application to a large dataset of volatilities demonstrate the competitive performance of the proposed method.
11footnotetext: Department of Economics, Università di Bologna. Email: matteo.barigozzi@unibo.it.22footnotetext: School of Mathematics, University of Bristol. Email: haeran.cho@bristol.ac.uk.33footnotetext: Department of Economics and Management, Universita’ di Pavia and School of Business, University of Leicester. Email: lorenzo.l.trapani@gmail.com.
@@ 1 Introduction
Factor models are, arguably, one of the most frequently employed tools to model and carry out inference when large-dimensional, vector-valued time series are available. Whilst a comprehensive review of the literature goes well above and beyond the scope of this paper, it is worth noting that factor models have a long history, dating back at least to the seminal paper by Spearman, (1904). Since being popularised by the contribution by Chamberlain and Rothschild, (1983), and since the development of the asymptotic theory to analyse large dimensional factor models (Bai,, 2003), their usage has become de rigueur in social sciences and economics where they have been applied to diverse fields such as business cycle analysis, asset pricing and economic monitoring and forecasting (see, inter alia, the review by Stock and Watson, (2011) for a comprehensive list of references).
@@
Where factor models are concerned, on the one hand, the analysis in Stock and Watson, (2009) and Bates et al., (2013) shows that, when changes are “sufficiently small” and do not involve a change in the number of common factors, the presence of change points is inconsequential for inference on the factor spaces. On the other hand, in several applications, there is no guarantee that the factor structure only undergoes small changes and that the number of common factors is time-invariant. Hence, it is not surprising that the literature has developed several methods to test, retrospectively, whether there is a change point or not. Breitung and Eickmeier, (2011), Corradi and Swanson, (2014), Han and Inoue, (2015), Yamamoto and Tanaka, (2015), Baltagi et al., (2017), Cheng et al., (2016) and Bai et al., (2024) propose several tests to check if there is a break, based on the idea that a change point in the loadings is observationally equivalent to a change in the covariance matrix of the common factors. Thus assuming homoscedastic factors, a change in the covariance matrix of the estimated common factors can only be due to a break in the loadings. See also Barigozzi and Trapani, (2020) and He et al., (2024) for alternative formulation of change point tests in an online context. Whilst the literature on change point testing is well-established, there is relatively little work on detection and estimation of (possibly) multiple change points in the factor structure with some exceptions (Barigozzi et al.,, 2018; Li et al.,, 2023; Bai et al.,, 2024).
@@
In this paper, we make an advance on the change point literature by proposing a method to estimate the number and locations of (possibly) multiple change points under a factor model. Specifically, we propose a procedure based on the so-called MOving SUMs (MOSUM) process which, since the seminal contribution by Hušková and Slabỳ, (2001), has been shown to have desirable properties in multiple change point estimation under weak assumptions on the distribution of the data. Versatility of the MOSUM procedure has been demonstrated in the context of detecting changes in the mean (Hušková and Slabỳ,, 2001; Eichinger and Kirch,, 2018), trend (Kim et al.,, 2024) and the drift of stochastic processes (Kirch and Klein,, 2024); see also Kirch and Reckruehm, (2024) for a general framework for multivariate time series segmentation based on an estimating function. Some recent contributions extend the use of this methodology to high-dimensional time series (Cho and Owens,, 2024; Cho et al.,, 2024), demonstrating its scalability and suggesting that it is worth exploring the performance of a MOSUM procedure in the context of large factor models.
@@
We build on that, as discussed above, a change in the loadings is observationally equivalent to a change in the covariance matrix of the common factors, and construct a MOSUM statistic based on the (moving) partial sums of the outer products of the estimated factors. In this respect, our statistic is akin to the one based on the (maximally selected) sequence of likelihood ratio tests proposed e.g. in Duan et al., (2023), save for the fact that it is based on using moving, as opposed to cumulative, partial sums. Upon testing for changes in the factor model using the MOSUM process, if the null hypothesis of no change is rejected, we perform multiple change point detection by estimating both the total number and the locations of the breaks. We establish the asymptotic null distribution of the MOSUM process as well as the consistency of the proposed change point detection procedure in estimating the total number and locations of the change points, with the accompanying rate of estimation for individual breaks. In doing so, we derive a Strong Invariance Principle for the MOSUM process based on the outer products of the estimated factors, which may be of independent interest.
@@
For a random variable $X$, we have $|X|_{\nu}=(\mathsf{E}(|X|^{\nu}))^{1/\nu}$. We denote: weak convergence on the space $\mathcal{D}[0,1]$ with $\overset{\mathcal{D}}{\rightarrow}$; =$\overset{\mathcal{D}}{=}$ is equality in distribution; $\rightarrow$ is the ordinary limit; $\Gamma(\cdot)$ is the Gamma function; $\log(x)$ is the natural logarithm of $x>0$; $\mathbb{I}_{\mathcal{A}}$ is an indicator function satisfying $\mathbb{I}_{\mathcal{A}}=1$ if the event $\mathcal{A}$ is true and $\mathbb{I}_{\mathcal{A}}=0$ otherwise. By $\mathbf{I}$, $\mathbf{O}$ and $\mathbf{0}$, we denote an identity matrix, a matrix of zeros and a vector of zeros whose dimensions depend on the context. For a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, we denote by $\mathbf{A}^{\top}$ its transpose and $\|\mathbf{A}\|$ its Euclidean norm, with $\Lambda_{\max}(\mathbf{A})$ and $\Lambda_{\min}(\mathbf{A})$ denoting its largest and the smallest eigenvalues in modulus. When $m=n$, we denote by $\mathsf{Vech}(\mathbf{A})$ the vector of length $m(m+1)/2$ that stacks the elements on and below the main diagonal of $\mathbf{A}$, and $\mathbf{L}_{m}$ the $m(m+1)/2\times m^{2}$-matrix satisfying $\mathsf{Vech}(\mathbf{A})=\mathbf{L}_{m}\mathsf{Vec}(\mathbf{A})$, where $\mathsf{Vec}(\cdot)$ is an operator that stacks the columns of the matrix into a vector. Conversely, we define $\mathbf{K}_{m}\in\mathbb{R}^{m^{2}\times m(m+1)/2}$ satisfying $\mathsf{Vec}(\mathbf{A})=\mathbf{K}_{m}\mathsf{Vech}(\mathbf{A})$, see Appendix A.12 of Lütkepohl, (2005). Finally, given two sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}=O(b_{n})$ if, for some finite positive constant $C$ there exists $N\in\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$ such that $|a_{n}||b_{n}|^{-1}\leq C$ for all $n\geq N$.
@@ 2.1 Model
Here, the matrix of loadings $\bm{\Lambda}_{j}=[\bm{\lambda}_{j,1},\ldots,\bm{\lambda}_{j,N}]^{\top}\in\mathbb{R}^{N\times r}$ has column rank $r_{j}\leq r$ fixed for all $N$, and “loads” the vector of random factors $\mathbf{f}_{t}=(f_{1,t},\ldots,f_{r,t})^{\top}$ onto the cross-sections of $\mathbf{X}_{t}$, and $\mathbf{e}_{t}=(e_{1,t},\ldots,e_{N,t})^{\top}$ denotes the idiosyncratic component. The model in (1) allows for the changes due to emergence or disappearance of factor(s) by permitting the ranks of $\bm{\Lambda}_{j}$ to vary, as well as rotational changes in the loading matrices. At the same time, the model is not identifiable in that the changes in the loadings are (observationally) equivalent to changes in the second-order properties of the common factor $\mathbf{f}_{t}$ – a fact, as mentioned in Introduction, frequently explored in the relevant literature for developing change point tests; see e.g. Bai et al., (2024). We further illustrate this point in the following example.
@@
Consider the factor model in (1) with $R=1$ and the single change point at $k_{1}=k^{\ast}$. For any $\bm{\Lambda}_{j}\in\mathbb{R}^{N\times r_{j}}$, $j=0,1$, we can find $\bm{\Lambda}\in\mathbb{R}^{N\times r}$ of full column rank with $r\geq\max(r_{0},r_{1})$ such that $\bm{\Lambda}_{j}=\bm{\Lambda}\mathbf{A}_{j}$ for $\mathbf{A}_{j}\in\mathbb{R}^{r\times r_{j}}$ of rank $r_{j}$. Then, we can re-write (1) as
@@
$\mathbf{X}_{t}=\bm{\Lambda}\sum_{j=0}^{R}\mathbf{A}_{j}\mathbf{f}_{t}\cdot\mathbb{I}_{\{k_{j}<t\leq k_{j+1}\}}+\mathbf{e}_{t}=:\bm{\Lambda}\mathbf{g}_{t}+\mathbf{e}_{t}$ (2)
@@
for $1\leq j\leq R$. Then, the problem of detecting the multiple change points in the loadings under (2), becomes that of detecting change points in the covariance of the pseudo factors $\{\mathbf{g}_{t}\}$.
@@ 2.2 Methodology
Let us suppose that the number of pseudo factors $r$ is known. Then, we can estimate the pseudo factors $\mathbf{g}_{t}$ as, up to an invertible transformation, $\sqrt{T}$ times the $r$ leading eigenvectors of the $T\times T$ matrix $(NT)^{-1}\mathbf{X}\mathbf{X}^{\top}$ with $\mathbf{X}=\mathbf{X}_{0:T}$. Denoting such estimator by $\widehat{\mathbf{g}}_{t}$, we propose to test for a change under the model in (1) and, if any, estimate the multiple change points, by scanning the data using the following MOSUM statistic:
@@
For the implementation of the MOSUM procedure, we require an estimator of the long-run covariance matrix $\mathbf{V}_{k}$. While we allow it to be location-dependent to account for the heteroscedasticity in the presence of change points, estimating a long-run covariance matrix of multivariate time series is well-known to be highly challenging. In the current setting, this is augmented by that the computation of $\mathcal{T}_{N,T,\gamma}(k)$ calls for the inverse of $\mathbf{V}_{k}$, which may bring further numerical instabilities. Therefore, we opt to use the following HAC-type estimator in place of $\mathbf{V}_{k}$,
@@
with some bandwidth $m\geq 1$. We later show that the matrix $\widehat{\mathbf{V}}$ provides a consistent estimator of the long-run covariance matrix of $\mathbf{M}_{N,T,\gamma}(k)$ under $\mathcal{H}_{0}$ (see Proposition 2). The Bartlett kernel in (5) is only one of the many possible choices, and we refer to Bai et al., (2024) for a comprehensive analysis of various kernel-based estimators of $\mathbf{V}$. For consistency of multiple change point detection, we only require that a positive definite matrix is used in place of $\mathbf{V}_{k}$ (Theorem 3). By default, we propose to adopt $\widehat{\mathbf{V}}$ (or its diagonal entries) which is shown to work well; see Section 4 for further discussions.
@@ 3.1 Assumptions
Since the seminal works by Wu, (2005) and Berkes et al., (2011) (see also Hörmann,, 2009), decomposable Bernoulli shifts have proven a very convenient way to model and study dependent time series, mainly due to their generality and since it is much easier to verify whether a sequence forms a decomposable Bernoulli shift than, e.g. verifying mixing conditions. Virtually all of the most commonly employed models in econometrics and statistics can be shown to generate decomposable Bernoulli shifts, such as ARMA and (G)ARCH processes, non-linear time series models (e.g. random coefficient autoregressive models and threshold models), Volterra series and data generated by dynamical systems; see Berkes et al., (2011), Aue et al., (2009) and Liu and Lin, (2009).
@@
    we suppose that $\mathbf{D}$ is invertible.
@@ 3.2 Asymptotic null distribution
The limiting law in Theorem 1 is analogous to those derived in Hušková and Slabỳ, (2001) and Eichinger and Kirch, (2018), modulo the fact that here, we deal with $d$-variate vectors and therefore the function $b_{d}(\cdot)$ depends on $d$. In contrast with a “standard” multivariate time series application, however, in our result, the cross-sectional dimension $N$ also plays a role through the definition of $\zeta$ in (7), which enters in the condition (8) made on the bandwidth $\gamma$. As a by-product, we establish the SIP of the process $\mathsf{Vech}(\widehat{\mathbf{g}}_{t}\widehat{\mathbf{g}}_{t}^{\top})$ after an appropriate centering (see Lemma A.11 in Appendix A.1), which proves crucial in deriving the asymptotic null distribution in (10).
@@
The condition in (8) places both upper and lower bounds on the bandwidth $\gamma$. Specifically, $\gamma$ is required to grow sufficiently faster than $T^{2\zeta}$ while satisfying $T^{-1}\gamma\rightarrow 0$, and the former restriction calls for larger $\gamma$ when $\mathbf{f}_{t}$ has fewer moments or when $N$ is small. We note that $\beta=\log(N)/\log(T)$ is known and does not need to be estimated. Conversely, $\nu$ is in general not known. We may select $\gamma$ satisfying (8) by plugging in an estimate of $\nu$, say $\widehat{\nu}$; alternatively, one can decide a value of $\nu$, say $\nu^{\ast}$, and test whether $\left|\mathbf{g}_{t}\right|_{\nu^{\ast}}<\infty$. In both cases, one difficulty is that $\mathbf{g}_{t}$ is not observable; however, deriving $\nu$ from the data $\mathbf{X}_{t}$ yields a lower bound.